OpenStudy (anonymous):

a spinner is divided into eight equal sections. five sections are red and three are green. if the spinner is spun three times, what is the probability that it lands on red exactly twice.

OpenStudy (anonymous):

75/512

OpenStudy (anonymous):

please show all work how did u get to that answer

OpenStudy (anonymous):

alright well you have 8 sections correct? when doing independent events you multiply each probability for each spin if you are trying to hit a red spot and there is 5 red spots that means the probability to hit a red spot is 5/8 so if you have to hit it twice you'll multiply thta probability twice lastly it says you can only exactly hit the red spots twice meaning your final spin must be green so the probability is 3/8 multiplying these all together \[\frac{5}{8}*\frac{5}{8}*\frac{3}{8}=\frac{75}{512}\]

OpenStudy (anonymous):

okay thank you very much

OpenStudy (anonymous):

that was one of the questions on the regents. the answer is not 75/512. the answer is 225/512 because 3C2(5/8)^2(3/8)^1 = 225/512